OpenStudy (dumbsearch2):

I would like to know, step by step, how to factor this polynomial: 24x^2+6x-30

OpenStudy (anonymous):

what is the greatest common factor of the coefficients (24,6,30)

OpenStudy (dumbsearch2):

?

OpenStudy (dumbsearch2):

GCF ?

OpenStudy (anonymous):

its 6. 6*4 = 24; 6*1 = 6; 6*5 = 30

OpenStudy (dumbsearch2):

how is GCF relevant to solving this polynomial ?

OpenStudy (anonymous):

you are FACTORING

OpenStudy (anonymous):

24x^2+6x-30 6(4x^2+x-5)

OpenStudy (dumbsearch2):

im confused my book says something different for thes type of problems "FIND TWO NUMBERS WHO PRODUCTS IS C AND WHOSE DIFFERENCE IS B AND

OpenStudy (anonymous):

the first thing you always do is factor out the greatest common factor. then you can start doing what your book says.

OpenStudy (dumbsearch2):

they don't say that

OpenStudy (dumbsearch2):

btw its trinomial not bynomial